What is the solution to the system of equations
x+3y=12
4x-y=-17

Answer by checkley71(8403)

X+3Y=12 NOW MULTIPLY THE SECOND EQUATION BY 3 & ADD THEM
12X-3Y=-51
------------------------------------------------------
13X=-39
X=-39/13
X=-3 ANSWER.
-3+3Y=12
3Y=12+3
3Y=15
Y=15/3
Y=5 ANSWER.
PROOF
4*-3-5=-17
-12-5=-17
-17=-17